OpenStudy (grace111):

Please help! Will reward ::: medal + fan + testimonial. N stands for number . Complete this table of values. (below)

OpenStudy (grace111):



OpenStudy (anonymous):

n=20 5n-2=98

OpenStudy (anonymous):

5n-2=38 38+2=40, 5n=40 40/5=8 n=8 in the bottom one

OpenStudy (anonymous):



OpenStudy (grace111):

thank you so much !

OpenStudy (anonymous):

your welcome